OpenStudy (anonymous):

in a flock of black and white sheep, 2 out of 5 sheep are white. if there are 8 more black sheep than white sheep, how many sheep are in the flock?

OpenStudy (anonymous):

as per the question for 1 more black sheep there are 2 white sheep thus the no. of white sheep=2*8=16 no of black sheep =2*8+8=24 thus the required no of sheep=16+24=40

OpenStudy (hba):

Ok I'll explain this as simply as possible- So to set this up as an equation, lets say the white sheep are x. Because this is a ratio problem, you need to set it up as a fraction. So because there are 8 more black sheep than there are white sheep, the radical will be x+8. Now the denominator has to be the total number of sheep, which will be x+8, the black sheep, +x, the white sheep, or 2x+8. Now your fraction will look like this: x+8 ----- 2x+8 Now you have to set this equal to the ratio of black sheep to the total number of sheep, which is 3:5 because the ratio of white sheep to the total number of sheep is 2:5. So now your equation will look like this: x+8 3 ------ = ----- 2x+8 5 So now you just have to cross multiply: 5(x+8)=3(2x+8) 5x+40=6x+24 -5x-24 -5x -24 16=x, so there are 16 white sheep. 16+8=24, so there are 24 black sheep, giving you a total of 40 sheep.

OpenStudy (anonymous):

thanks guyssssss